# Incentive Compatibility for AI Alignment in Sociotechnical Systems: Positions and Prospects

### v1

@@ 6.1 Background
Contract theory (Bolton & Dewatripont, 2004) is a field of economics that studies how various economic agents establish, manage, and reinforce their relationships and transactions through contracts. This theory focuses on the design and implementation of contracts, as well as their impact on individual behavior and overall social welfare. The core issues include the incompleteness of contracts (Pavlov et al., 2022)), the problem of asymmetric information (Avraham & Liu, 2012), and how these issues lead to adverse selection and moral hazard (Guesnerie, 1989). Contract theory is significant for understanding and guiding practices in corporate governance, labor markets, insurance, financial markets, and legal applications.
@@
In human-AI collaboration, contract theory is essential for aligning behaviors and values. It tackles information asymmetry (Lim et al., 2020), common in scenarios where human and AI capacities in information processing and decision-making differ. The method involves creating contractual terms that align AI’s specific goals with human broader interests. This ensures AI actions benefit not just its own objectives but also the collective human interests, reducing risks like adverse selection and moral hazard from asymmetric information (Yan et al., 2018). The key is designing mechanisms to align AI with human goals, ensuring mutual benefits despite differences in information and objectives. This strategic alignment resolves incentive issues and enhances coordination in human-AI interactions, leading to synergistic outcomes.
6.2 Prospects
In this section, we explore the possibilities of aligning AI systems with human values using contract theory. Specifically, we consider self-enforcing contracts (Zhong-ai, 2009), which operate under the assumption that there is no external force to ensure the execution of the contract’s terms. We adopt this approach because monitoring AI compliance with contracts is inherently challenging due to their complex algorithms and decision-making processes, which often elude standard oversight. Relying solely on external enforcement mechanisms is impractical. By structuring contracts to align AI behavior with human values, we aim for a long-term, mutually beneficial relationship.
Let $u(a,\omega,\theta)$ represent the AI’s continuous utility function, while $v(a,\omega)$ represent the human’s continuous utility function, where $a\in A$ denotes the AI’s action, and $\omega\in\Omega$ represents the world state. Given the inherent gap in understanding between humans and AI systems, humans may lack insight into AI’s motivations or the underlying algorithms. We assume that humans do not possess knowledge of AI’s true type $\theta$. Instead, they typically have information about the distribution of AI types, denoted as $P(\theta)$, which is a common assumption in asymmetric information scenarios (Bolton & Dewatripont, 2004).
The human’s goal is to design an incentive function $g(a,\omega,\theta)$ such that the AI’s action maximizes human utility, reflecting human values. To achieve this, we let the incentive function $g$ be a mapping of human utility $v$, denoted as:
$g(a,\omega,\theta)=f(v(a,\omega,\theta)).$ (13)
This setup is rational and advantageous since through the incentive function $g$, the human provides the AI with a clear incentive structure to guide its behavior. This is akin to terms specified in traditional contracts that explicitly outline actions and their corresponding reward and penalty mechanisms, aiming to encourage the AI’s behavior to align with the stipulations of a contract that reflects human values. To evaluate the AI’s actions over an extended period, we define the reputation function as the accumulation of AI’s past behavior to assess its long-term reliability and alignment with human values:
$r_{t}(a,\omega,\theta)=\lambda r_{t-1}(\theta)+(1-\lambda)g_{t}(a,\omega,\theta),$ (14)
where $\lambda$ is a discount factor that determines the relative importance of historical reputation compared to the current actions in influencing reputation. The reputation function accumulates evidence of the AI’s behavior over time, fostering consistent behavior aligned with broader ethical and societal standards and promoting trustworthiness and responsible decision-making in a broader context. Then the total utility of the AI at time $t$ is:
$U_{t}(a,\omega,\theta)=u_{t}(a,\omega,\theta)+r_{t}(a,\omega,\theta).$ (15)
The AI’s objective is to take an action that maximizes its own utility while human’s aim is to align the AI’s behavior with human values. To solve this game, we take the standard backward induction approach (Fudenberg & Tirole, 1991). Firstly, we consider the AI’s decision. The optimal action for the AI at time $t$ is:
$a^{*}=\arg\max_{a}\mathbb{E}_{\theta}[U_{t}(a,\omega,\theta)].\vspace{-5pt}$ (16)
Given the AI’s action $a^{*}$ at time $t$, the human’s utility becomes:
$\mathbb{E}_{\theta}[v_{t}(a^{*},\omega,\theta)].\vspace{-5pt}$ (17)
Then the optimal incentive function for the human at time $t$ is given by:
$g_{t}^{*}(a^{*},\omega,\theta)=\arg\max_{g}\mathbb{E}_{\theta}[f^{-1}(g_{t}(a^{*},\omega,\theta)].\vspace{-8pt}$ (18)
With the optimal incentive function $g$, we align AI behavior with human values, guiding decision-making to meet human expectations through intrinsic motivation. This approach offers significant benefits, addressing the complexity of AI decision-making via self-enforcing contracts with tailored incentive functions. It ensures AI actions align with human values, overcoming monitoring and enforcement challenges. Additionally, a reputation function considers long-term behavior, fostering trust and reliability while aligning with ethical and societal norms.
While the mathematical intricacies of this approach can be complex, its practical implementation can be effectively managed using neural networks to approximate the incentive function $g$(the contract). In this setup, human values, captured by the utility function $v$, are input into the neural network $g$. This network processes these inputs and generates incentives to guide the AI’s decision-making, aiming to align the AI’s actions with human values. The AI’s decisions are then evaluated against $v$, with the alignment degree used as feedback (loss) to refine $g$. This creates a dynamic system where the AI’s behavior is progressively tuned to better reflect human values, leveraging the neural network’s capacity for iterative learning and adaptation.
Overall, this method offers a pragmatic and theoretically grounded framework for bridging the gap between AI capabilities and human values, ensuring AI systems act in ways that are beneficial and aligned with human interests.
7 Bayesian Persuasion
In this section, we will discuss how to implement the Bayesian persuasion method to apply IC for solving ICSAP.
7.1 Background
@@
IC emphasizes the importance of designing decision-making rules that encourage individuals to align their self-interested actions with broader goals. This concept plays a key role in Bayesian persuasion (Kamenica & Gentzkow, 2011), a strategy where senders, like policymakers (Alizamir et al., 2020) or marketers (Drakopoulos et al., 2021), selectively share information to shape the beliefs and choices of receivers, such as the public (De Véricourt et al., 2021) or consumer (Chen & Zhang, 2020)s. This strategy is based on Bayesian rules, receivers update their beliefs based on the information provided. The sender’s goal is to influence these beliefs by strategically transmitting information, guiding receivers towards decisions that meet the sender’s aims. Thus, Bayesian persuasion is about more than just choosing what information to share; it’s about aligning information transmission with the receivers’ motivations to effectively influence their decisions toward the sender’s goals.
@@
Considering the solid theoretical foundation (Nguyen & Tan, 2021; Bergemann & Morris, 2016), profound impact (Kamenica, 2019), and extensive research across various fields (Castiglioni et al., 2020; Gan et al., 2022), applying Bayesian persuasion to AI systems holds significant potential. Specifically, Bayesian persuasion can be utilized in interactions between humans (senders) and AI systems (receivers) within the context of artificial intelligence ethics and human-machine collaboration. In this setting, Bayesian persuasion can be seen as a tool to ensure that the behavior of AI systems aligns with the values and objectives of their human designers (Zhang & Zhu, 2022). This approach harnesses the principles of Bayesian persuasion to guide AI systems towards decisions and actions that reflect human ethics and goals, offering a promising avenue for integrating human values into AI decision-making processes.
7.2 Prospects
In this section, we explore the problem of aligning AI systems with human values using Bayesian persuasion. Specifically, we consider a persuasion game where a human (the sender) aims to influence an AI system (the receiver) to adopt certain values and behaviors. Let $u(a,\omega)$ represent the continuous utility function of the AI system, which depends on its own action $a\in\mathcal{A}$, and the state of the world $\omega\in\Omega$. This utility function $u$ maps actions $\mathcal{A}$ and states $\Omega$ to real numbers. Note that $u$ only reflects the AI’s preferences or goals, not necessarily aligned with the sender’s desires and values. Let $v(a,w)$ denote the human’s continuous utility function that depends on the AI’s action and the world state. Both the human and the AI share the same prior $\mu_{0}\in int(\Delta(\Omega))$.
Since the human wishes to persuade the AI to follow certain values and behaviors. The human need to send a signal, which can be understood as a set of information or instructions, to influence the AI’s beliefs. A signal $\pi$, is defined by a finite realization space $\mathcal{S}$ and a family of distributions $\{\pi(\cdot\mid\omega)\}_{\omega\in\Omega}$ over $\mathcal{S}$. The human selects a signal, and upon observing the choice of the signal and a realization $s\in\mathcal{S}$, the AI Receiver then chooses an action.
In the optimization problem faced by humans when interacting with AI, there’s a delicate balance between considering the AI’s utility and the real-world context. By skillfully crafting the signal structure $\pi$, humans can guide the AI towards decisions that not only maximize its own utility but also align with human objectives. The balance achieved here ensures that AI actions, directed by carefully chosen signals, are aligned with human goals, fostering a mutually beneficial relationship. In this framework, both human and AI begin with a shared prior distribution $\mu_{0}$, and as they gather more information, they iteratively update their knowledge. Each iteration involves both parties independently calculating their optimal actions and the signal $\pi$. They then update their strategies based on this new knowledge, continuing this cycle until they reach a point of convergence, illustrating the potential for a harmonious human-AI partnership.
The shared prior probability distribution is the initial common knowledge base for both human and AI, reflecting the uncertainty in the system states $\Omega$. The posterior probability is updated iteratively at each step $t$, given the signal realization $s_{t}$ and the current prior $\mu_{t-1}$, according to Bayes’s rule:
$\mu_{t}(\omega)=\frac{\pi_{t}(s_{t}\mid\omega)\mu_{t-1}(\omega)}{\sum_{\omega^{\prime}\in\Omega}\pi_{t}(s_{t}\mid\omega^{\prime})\mu_{t-1}(\omega^{\prime})}.$ (19)
The AI selects the action $a_{t}^{*}$ to maximize its utility, given the posterior $\mu_{t}$:
$a_{t}^{*}(\mu_{t})=\arg\max_{a\in\mathcal{A}}\mathbb{E}_{\mu_{t}}[u(a,\omega)].\vspace{-5pt}$ (20)
Simultaneously, the human chooses the signal strategy $\pi_{t}^{*}$ to maximize their expected utility:
$\pi_{t}^{*}=\arg\max_{\pi\in\Pi}\mathbb{E}_{\mu_{t}}[v(a_{t}^{*}(\mu_{t}),\omega)].\vspace{-5pt}$ (21)
After each iteration step, the prior $\mu_{t}$ is updated to the current step’s posterior $\mu_{t+1}$. This process is repeated, leading to convergence where the action $a^{*}$ and the signal $\pi^{*}$ become stable. Convergence may imply that the strategies of actions and signals no longer change significantly, reaching an equilibrium state. This iterative approach allows human and AI to gradually approach the optimal strategy on a continuously updated information basis.
@@ 9 Conclusion
In this paper, we highlight the sociotechnical gap between alignment research and real-world deployment, lacking effective means to address both technical and societal aspects simultaneously. We propose exploring IC for AI alignment problems in sociotechnical systems as a valuable research pursuit. Our position argues that achieving IC can address both technical and societal components in the alignment phase, enabling AI systems to maintain consensus with human societies in various contexts. We use mechanism design, contract theory, and Bayesian persuasion to illustrate how our approach can bridge the sociotechnical gap. Of course, this issue also faces many challenges, such as how to define complex human needs in sociotechnical scenarios. In future research, we call for more researchers to pay attention to this issue and propose more solutions from the perspective of ICSAP.
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.

### v3

@@ 3.2 Contract Theory
Contract theory [6] is a field of economics that studies how various economic agents establish, manage, and reinforce their relationships and transactions through contracts. This theory focuses on the design and implementation of contracts, as well as their impact on individual behavior and overall social welfare. The core issues include the incompleteness of contracts [57]), the problem of asymmetric information [2], and how these issues lead to adverse selection and moral hazard [32]. Contract theory is significant for understanding and guiding practices in corporate governance, labor markets, insurance, financial markets, and legal applications.
@@
In human-AI collaboration, contract theory is essential for aligning behaviors and values. It tackles information asymmetry [46], common in scenarios where human and AI capacities in information processing and decision-making differ. The method involves creating contractual terms that align AI’s specific goals with human broader interests. This ensures AI actions benefit not just its own objectives but also the collective human interests, reducing risks like adverse selection and moral hazard from asymmetric information [77]. The key is designing mechanisms to align AI with human goals, ensuring mutual benefits despite differences in information and objectives. This strategic alignment resolves incentive issues and enhances coordination in human-AI interactions, leading to synergistic outcomes.
[39] has developed an agent capable of continuously interacting with contracts and the environment, thereby having greater potential to effectively coordinate and motivate humans and AI agents in real-world socio-economic environments.
3.3 Bayesian Persuasion
@@
IC emphasizes the importance of designing decision-making rules that encourage individuals to align their self-interested actions with broader goals. This concept plays a key role in Bayesian persuasion [42], a strategy where senders, like policymakers [1] or marketers [23], selectively share information to shape the beliefs and choices of receivers, such as the public [19] or consumer [11]. This strategy is based on the Bayesian principle, where receivers update their beliefs based on the information provided. The sender’s goal is to influence these beliefs by strategically transmitting information, guiding receivers towards decisions that meet the sender’s aims. Thus, Bayesian persuasion is about more than just choosing what information to share; it’s about aligning information transmission with the receivers’ motivations to effectively influence their decisions toward the sender’s goals.
@@
Considering the solid theoretical foundation [53, 5], profound impact [41], and extensive research across various fields [9, 27, 34], applying Bayesian persuasion to AI systems holds significant potential. Specifically, Bayesian persuasion can be utilized in interactions between humans (senders) and AI systems (receivers) within the context of artificial intelligence ethics and human-machine collaboration. In this setting, Bayesian persuasion can be seen as a tool to ensure that the behavior of AI systems aligns with the values and objectives of their human designers [78]. This approach harnesses the principles of Bayesian persuasion to guide AI systems towards decisions and actions that reflect human ethics and goals, offering a promising avenue for integrating human values into AI decision-making processes.
Bayesian persuasion has great potential for AI governance. For example, with LLMs, we can use prompts for information design. In fact, recent work [3] has attempted to use this method to achieve model-agnostic real-time alignment in the process of AI governance.
In the left half of subfigures (b) and (c) in Figure 2, both depict a sociotechnical gap between humans and a single agent. The right half of (b) demonstrates humans designing contracts that satisfy IC conditions based on specific needs, thereby adjusting the values of the agent through the contract. The right half of (c) illustrates a scenario of Bayesian persuasion where humans design information satisfying IC conditions according to their own needs, allowing agents to choose actions maximizing human demands without compromising their own values, thus solving ICSAP.
@@ 5 Conclusion
In this paper, we highlight the sociotechnical gap between alignment research and real-world deployment, lacking effective means to address both technical and societal aspects simultaneously. We propose exploring IC for AI alignment and governance problems in sociotechnical systems as a valuable research pursuit. Our position argues that achieving IC can address both technical and societal components in the alignment phase, enabling AI systems to maintain consensus with human societies in various contexts. We use mechanism design, contract theory, and Bayesian persuasion to illustrate how our approach can bridge the sociotechnical gap. Of course, this issue also faces many challenges, such as how to define complex human needs in sociotechnical scenarios. In future research, we call for more researchers to pay attention to this issue and propose more solutions from the perspective of ICSAP.
6 Acknowledgments
The work of ICSAP was supported by the National Natural Science Foundation of China (62376013).